What is the change in length of a 30.0-cm-long column of mercury if its temperature changes from 37.0ºC to 40.0ºC , assuming the mercury is constrained in a tube so that the volume only changes in one dimension? (answer in ×10−6 m
Answer #1

gamma = 181*10^-6

initial volume = V1 = A*L1

final volume = V2 = A*L2

here area A remains same


gamma = dV/(v1*dT)

gamma = A*dL/dT


181*10^-6 = dL/(30*(40-37))


dL = 16290*10^-6 m